What is the slope of any line perpendicular to the line through points (-2,3) and (4,5)?
Found 2 solutions by Alan3354, london maths tutor:
Answer by Alan3354(69443)

Given 2 points example.
Find the equation of the line thru the points (2,1) and (3,5)
---------------------
slope, m = diffy/diffx
m = (5-1)/(3-2)
m = 4
----------------
The slope of lines perpendicular is the negative inverse.
m = -1/4

Answer by london maths tutor(243)

The slope of any line perpendicular to the line = -1/m
where m is the gradient of the line.
Formular for finding gradient using two points = [Y,2 - Y,1]/[X,2-X,1]
Gradient = [5-3]/[4-(-2)]
= 2/6
=1/3
Therefore, the slope of the prependicular line = -3